# Testing the Froggatt-Nielsen Mechanism with Lepton Violation

### v1

@@ I Introduction
Standard Model (SM) fermions exhibit a broad range of masses and mixing angles with distinct patterns. While technically natural, this dramatic variation invites an explanation beyond the SM. The Froggatt-Nielsen (FN) mechanism [1] provides an elegant and economical framework for explaining this structure through a spontaneously broken horizontal symmetry.
@@
While this mechanism has been extensively studied in the quark sector, its application to leptons introduces new challenges [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Unlike quarks, whose masses and mixings all arise from Dirac-like Yukawa couplings, the underlying mechanism for generating neutrino masses is currently unknown, leading to a greater variety of possible implementations. Furthermore, the Pontecorvo Maki Nakagawa Sakata (PMNS) matrix does not exhibit any hierarchies and there are large experimental uncertainties on the parameters of this sector; indeed, the precise values of neutrino masses are currently unknown.
In this work we address these challenges systematically for the first time through a two step procedure:
  1. 1.
    Identify realistic natural FN textures
@@
    We first scan over a wide range of FN charge assignments (“textures”) for the leptons, and identify hundreds of combinations that naturally yield viable lepton masses and mixings. In our treatment, all other parameters (e.g. the coefficients of FN-preserving operators) are chosen to be generic order-one numbers such that the hierarchies in the lepton sector arise entirely from the charge assignments and corresponding FN spurion insertions.
  2. 2.
@@
    Predict CLFV and $0\nu\beta\beta$ for these textures Although the SM does not predict observable levels of charged lepton flavor violation (CFLV) or neutrinoless double beta decay ($0\nu\beta\beta$), such processes can be greatly enhanced in the presence of higher-dimensional operators that arise from integrating out the FN sector. The relative importance of different observable processes is dictated by texture-specific selection rules for each model. We therefore calculate the predicted rates of lepton violating processes for the identified realistic textures within the Standard Model Effective Theory (SMEFT) framework, which allows us to evaluate the experimental prospects of the leptonic FN mechanism as a whole.
We find that Dirac and Majorana FN models predict characteristic CLFV signatures distinct from anarchic scenarios, with muon processes such as $\mu\to e\gamma$, $\mu\to 3e$, and muon-to-electron conversion emerging as the most sensitive probes, and correlations between measurements revealing details of the flavor charges. Interestingly, type-I seesaw FN models are beyond the foreseeable reach of CLFV probes, and purely leptonic FN models of any type only rarely generate observable flavor-violating signals at proposed muon colliders, though a future analysis that includes a FN quark sector could reveal motivated signals. However, both Majorana and type-I seesaw models do make distinct predictions for the rate of $0\nu\beta\beta$ decay. Finally, a future measurement of $\sum m_{\nu}$ that indicates inverted ordering of the neutrino masses would strongly disfavor Dirac FN models. Overall, our analysis reveals how a variety of low-energy, high-energy, and cosmological probes can provide a direct window into the dynamics underlying lepton flavor.
This letter is organized as follows. Section II reviews the FN mechanism and its implementation within different neutrino mass generation mechanisms. Section III outlines our numerical approach for identifying realistic FN textures and their observables. Key results, including top textures and specific experimental predictions, are presented in Section IV, followed by conclusions in Section V. Details on the numerical methods used and supplemental discussion on models and phenomenology are given in the Appendices.
@@ II Froggatt-Nielsen Neutrino Masses
The FN mechanism introduces a $U(1)_{\rm FN}\,$  symmetry which is spontaneously broken at a UV scale $\Lambda$ by the vacuum expectation value (VEV) of a heavy scalar $\phi$.
$\epsilon\equiv\frac{\phi}{\Lambda}\ll 1$ (1)
is the spurion associated with the breaking of $U(1)_{\rm FN}\,$. Without loss of generality, we take the FN charge of $\phi$ (SM Higgs) to be $X_{\phi}=1$ ($X_{H}=0$), and the VEV of $\phi$ to be in the positive real direction. The SM Yukawa couplings only arise through $\epsilon$ insertions after FN breaking.
@@
where $L_{i}$ are the left-handed (LH) lepton doublets, $e_{j}$ the right-handed (RH) charged lepton singlets, $H$ is the Higgs doublet, $c^{\ell}$ a coupling matrix with $\mathcal{O}(1)$ entries, and $i,j=1,2,3$ label the fermion generations. The resulting Yukawa matrix is
$(Y_{\ell})_{ij}\equiv c^{\ell}_{ij}\epsilon^{n^{\ell}_{ij}}~{}~{},~{}~{}~{}n^{\ell}_{ij}\equiv|X_{L_{i}}+X_{\bar{e}_{j}}|~{}~{},$ (3)
naturally generating hierarchies when $\epsilon\ll 1$. Diagonalizing the Yukawas yields the charged lepton masses
$Y_{\ell}=U_{\ell}\hat{Y}_{\ell}W_{\ell}^{\dagger}~{}~{}\implies~{}~{}\hat{m}_{\ell}=\frac{v}{\sqrt{2}}\hat{Y}_{\ell}~{}$ (4)
after electroweak symmetry breaking (EWSB), where $\hat{Y}_{\ell}$ is diagonal, $U_{\ell}$ and $\ W_{\ell}$ are unitary, and $v=246$ GeV.
To account for neutrino masses, we examine three generation mechanisms within the FN framework.
@@
${}~{}~{}\mathcal{L}_{D}\supset c^{\nu}_{ij}\epsilon^{n^{\nu}_{ij}}H{L}_{i}N_{j}~{}~{},~{}~{}n^{\nu}_{ij}\equiv|X_{L_{i}}+X_{N_{j}}|~{},\!\!\!\!\!$ (5)
where $c^{\nu}$ is a symmetric matrix with order-one entries. In analogy with charged leptons, the neutrino Yukawa matrix can then be written as $(Y_{\nu})_{ij}\equiv c^{\nu}_{ij}\epsilon^{n^{\nu}_{ij}}$ and yields
$Y_{\nu}=U_{\nu}\hat{Y}_{\nu}W_{\nu}^{\dagger}~{}\implies~{}\hat{m}_{\nu}=\frac{v}{\sqrt{2}}\hat{Y}_{\nu},$ (6)
where $\hat{Y}_{\nu}$ is the diagonal Yukawa matrix, $U_{\nu}$ and $W_{\nu}$ are unitary matrices, $\hat{m}_{\nu}$ is the diagonal mass matrix, and the PMNS matrix can be written as $V\equiv U^{\dagger}_{\ell}U_{\nu}$.
@@
Majorana – The Weinberg operator [30]
${}~{}~{}~{}~{}~{}~{}\mathcal{L}_{W}\!\!\supset\!-\frac{c_{ij}^{W}\epsilon^{n^{W}_{ij}}}{\Lambda_{W}}(L_{i}H)(L_{j}H),$ (7)
@@
Seesaw – We consider a type-I seesaw scenario [31, 32, 33, 34] where flavor breaking and lepton violation are governed by the FN mechanism. In this case, neutrinos receive both Dirac and Majorana mass terms:
@@
where $c^{\nu}$ and $c^{M}$ are matrices of order-one coefficients, $M$ is the Majorana mass scale of the RH neutrinos, and
$~{}~{}~{}n^{\nu}_{ij}\equiv|X_{L_{i}}+X_{N_{j}}|~{}~{},~{}~{}n^{M}_{ij}\equiv|X_{N_{i}}+X_{N_{j}}|~{}.$ (10)
In the seesaw limit, the Dirac contribution is small and the diagonal neutrino mass matrix is
$\hskip 25.6073pt\hat{m}_{\nu}\approx\frac{v^{2}}{2M}U_{\nu}^{T}\!\left(c^{\nu}\epsilon^{n^{\nu}}\right)\!\!\left(c^{M}\epsilon^{n^{M}}\right)^{-1}\!\left(c^{\nu}\epsilon^{n^{\nu}}\right)^{T}\!U_{\nu},~{}~{}~{}~{}~{}~{}$ (11)
@@
Note that most explicit models of neutrino mass generation can be accommodated within these scenarios. Notably, a type-II seesaw variation [35, 36, 37, 38, 39] within our framework yields a Weinberg operator whose suppression scale $\Lambda_{W}$ can be written in terms of $\epsilon$ and additional FN charges, but is otherwise identical (see Appendix C.3).
@@ III Methods
Identifying Realistic Natural FN Textures – We begin by identifying viable FN textures that adequately reproduce the observed lepton masses and mixing parameters for a common value of $\epsilon$ in Eq. (1). To ensure that, for a given texture, all hierarchical structure arises only from $\epsilon$, we demand that all other free parameters are order-one numbers. In our analysis, these free parameters are encoded in the coefficients
$~{}~{}c^{\ell}_{ij},~{}~{}c^{\nu}_{ij},~{}~{}c^{W}_{ij},~{}~{}c^{M}_{ij},~{}~{}$ (12)
from Eqs. (2), (5), (7), and (9), respectively. We scan over a wide range of possible FN charge assignments and determine the fraction of random order-one values in Eq. (12) that reproduce all observed masses and mixings within some tolerance.
@@
Specifically, we adopt and extend the Bayesian-inspired method of Ref. [29] to the lepton sector – see Appendix A for details. For each mass generation mechanism, we consider all textures with charges $|X|\leq 7$, and, for Dirac RH neutrinos, $|X|\leq 9$ as required to obtain viable models. For each texture, we generate random coefficient matrices $c_{ij}$, with each $\log_{10}c_{ij}$ sampled from a normal distribution centered on zero with standard deviation $\sigma=0.3$, and phases sampled uniformly over $[0,2\pi]$. (Alternative choices of reasonable “$\mathcal{O}(1)$” priors do not meaningfully affect results.)
For each choice of coefficients, we compute observables
$\mathcal{O}=\left\{m_{\ell}~{},~{}\Delta m^{2}_{ij}~{},~{}|V_{ij}|~{},~{}\sum m_{\nu}\right\},$ (13)
corresponding to the charged lepton masses, neutrino mass-squared differences, the absolute value of the PMNS matrix elements and the sum of neutrino masses. The fractional deviation of each FN prediction $\mathcal{O}_{\rm FN}$ is
$\delta_{\mathcal{O}}\sim\frac{\mathcal{O}_{\rm FN}}{\mathcal{O}_{\rm exp}},$ (14)
where $\mathcal{O}_{\rm exp}$ is the experimentally measured value. (For the more rigorous definition of $\delta_{\mathcal{O}}$ used in our scans, see Appendix A.1.) We maximize over all observables in Eq. (13) to obtain the overall experimental deviation:
$\delta_{\mathrm{max}}\equiv\max_{\mathcal{O}}\,(\delta_{\mathcal{O}}).$ (15)
Next, we adjust $\epsilon$ and, where applicable, $\Lambda_{W}$ or $\Lambda$, to minimize $\delta_{\mathrm{max}}$. For each texture, this process is repeated many times for many choices of order-one coefficients from Eq. (12). Textures that naturally resemble our world will have $\delta_{\mathrm{max}}\sim\mathcal{O}(1)$ for a large fraction coefficient choices. To compare textures, we define
$\mathcal{F}_{x}\equiv\text{\% of coeff. choices for which}~{}\delta_{\mathrm{max}}\leq x,$ (16)
allowing for textures to be ranked by $\mathcal{F}_{x}$ for different $x$.
@@
where $c_{ijkl}$ are ${\cal O}(1)$ coefficients and we have defined
$n_{ijkl}\equiv|X_{\psi_{i}}-X_{\psi_{j}}+X_{\psi_{k}}-X_{\psi_{l}}|.$ (18)
Note that we switched to four-component fermion notation to match SMEFT conventions, but we write $n_{ijkl}$ in terms of the FN charges of the corresponding 2-component fermion fields. For dipole operators, we have
@@
In our analysis, ${\cal O}(1)$ Wilson coefficients are generated in the FN basis at a high scale $\Lambda$ and subsequently rotated into the mass basis using the values of $\epsilon$ and the $U$ and $W$ matrices from Eqs. (4), (6), (8), (11), built using the coefficients in Eq. (12) appropriate to each scenario. We calculate low-energy observables, including two and three-body CLFV decays of muons and taus, as well as muon-to-electron conversion in nuclei, using the flavio package [40]. Matching to the relevant energy scales is performed with wilson [41] (running was found to be a negligible effect). Based on experimental constraints, we derive lower bounds on $\Lambda$. Finally, using these bounds, we estimate the highest possible rates of $e\mu$, $e\tau$, and $\mu\tau$ production at future $e^{+}e^{-}$ and $\mu^{+}\mu^{-}$ colliders. For comparison, we also compute the predictions for a fully anarchic or null texture, where all charges are set to zero, and the Wilson coefficients are assumed to be all $\mathcal{O}(1)$ at a fixed scale. Further details on our methodology can be found in Appendix A.3.

### v2

# Testing the Froggatt-Nielsen Mechanism with Lepton Flavor and Number Violating Processes

@@ I Introduction
Standard Model (SM) fermions exhibit a broad range of masses and mixing angles with distinct patterns. While technically natural, this dramatic variation invites an explanation beyond the SM. The Froggatt-Nielsen (FN) mechanism [Froggatt:1978nt] provides an elegant and economical framework for explaining this structure through a spontaneously broken horizontal symmetry.
@@
This mechanism has been studied in both the quark and lepton sector [Leurer:1992wg, Leurer:1993gy, Dudas:1996fe, Irges:1998ax, Sato:2000ff, Sato:2000kj, Suematsu:2001sp, Dreiner:2003hw, Nir:2004my, Kamikado:2008jx, Plentinger:2008up, Buchmuller:2011tm, Krippendorf:2015kta, Bauer:2016rxs, Ema:2016ops, Calibbi:2016hwq, Nishiwaki:2016xyp, Vien:2019zhs, Feruglio:2019ybq, Berger:2019rfk, Bordone:2019uzc, Smolkovic:2019jow, Fedele:2020fvh, Nishimura:2020nre, Aloni:2021wzk, Ringe:2022rjx, Asadi:2023ucx, Qiu:2023igq, Cornella:2023zme, Nishimura:2024apb, Ibe:2024cvi]. The lepton sector, which is the focus of this work, introduces the additional challenge of integrating neutrinos into the framework. Unlike quarks, whose masses and mixings all arise from Dirac-like Yukawa couplings, the underlying mechanism for generating neutrino masses is currently unknown, leading to a greater variety of possible implementations. Furthermore, the Pontecorvo Maki Nakagawa Sakata (PMNS) matrix does not exhibit any hierarchies and there are large experimental uncertainties on the parameters of this sector; indeed, the precise values of neutrino masses are currently unknown.
In this work we address these challenges systematically to present for the first time a global picture of the expected relative magnitudes of CLFV observables resulting from a large number of phenomenologically viable lepton FN textures. We also present the correlations between observables which could in principle be used to discriminate between textures. We achieve this through a two step procedure:
@@
1. Identify realistic natural FN textures: We first scan over a wide range of FN charge assignments (“textures”) for the leptons, and identify hundreds of combinations that naturally yield viable lepton masses and mixings. In our treatment, all other parameters (e.g. the coefficients of FN-preserving operators) are chosen to be generic order-one numbers such that the hierarchies in the lepton sector arise entirely from the charge assignments and corresponding FN spurion insertions.
@@
2. Predict CLFV and $0\nu\beta\beta$ for these textures: Although the SM does not predict observable levels of charged lepton flavor violation (CLFV) or neutrinoless double beta decay ($0\nu\beta\beta$), such processes can be greatly enhanced in the presence of higher-dimensional operators that arise from integrating out the FN sector. The relative importance of different observable processes is dictated by texture-specific selection rules for each model. We therefore calculate the predicted rates of lepton violating processes for the identified realistic textures within the Standard Model Effective Theory (SMEFT) framework, which allows us to evaluate the experimental prospects of the leptonic FN mechanism as a whole.
We consider neutrino mass generation by three mechanisms: Dirac, generic Majorana via a Weinberg operator, and type-I seesaw. [^1] Note that while many type-I seesaw scenarios can be mapped directly to a Weinberg operator with a single Weinberg scale, as in the Majorana case, if the RH neutrinos have flavor charge the type-I seesaw would effectively generate a Weinberg operator where the suppression scale has flavor structure. Our analysis reveals how a variety of low-energy, high-energy, and cosmological probes can provide a direct window into the dynamics underlying lepton flavor.
@@ II Froggatt-Nielsen Neutrino Masses
The FN mechanism introduces a $U(1)_{\rm FN}\,$  symmetry which is spontaneously broken at a UV scale $\Lambda$ by the vacuum expectation value (VEV) of a heavy scalar $\phi$. $\epsilon\equiv\phi/\Lambda\ll 1$ is the spurion associated with the breaking of $U(1)_{\rm FN}\,$. Without loss of generality, we take the FN charge of $\phi$ (SM Higgs) to be $X_{\phi}=1$ ($X_{H}=0$), and the VEV of $\phi$ to be in the positive real direction. The SM Yukawa couplings only arise through $\epsilon$ insertions after FN breaking.
@@
where $L_{i}$ are the left-handed (LH) lepton doublets, $e_{j}$ the right-handed (RH) charged lepton singlets, $H$ is the Higgs doublet, $c^{\ell}$ a coupling matrix with $\mathcal{O}(1)$ entries, and $i,j=1,2,3$ label the fermion generations. The resulting Yukawa matrix is $(Y_{\ell})_{ij}\equiv c^{\ell}_{ij}\epsilon^{n^{\ell}_{ij}},n^{\ell}_{ij}\equiv|X_{L_{i}}+X_{\bar{e}_{j}}|,$ naturally generating hierarchies when $\epsilon\ll 1$. Diagonalizing the Yukawas yields the charged lepton masses $Y_{\ell}=U_{\ell}\hat{Y}_{\ell}W_{\ell}^{\dagger}\implies\hat{m}_{\ell}=\frac{v}{\sqrt{2}}\hat{Y}_{\ell}~$ after electroweak symmetry breaking (EWSB), where $\hat{Y}_{\ell}$ is diagonal, $U_{\ell}$ and $W_{\ell}$ are unitary, and $v=246$ GeV. To account for neutrino masses, we examine three generation mechanisms within the FN framework.
@@
$\mathcal{L}_{D}\supset c^{\nu}_{ij}\epsilon^{n^{\nu}_{ij}}H{L}_{i}N_{j}~~,~~n^{\nu}_{ij}\equiv|X_{L_{i}}+X_{N_{j}}|~,\!\!\!\!\!$ (2)
where $c^{\nu}$ is a matrix with order-one entries. In analogy with charged leptons, the neutrino Yukawa matrix can then be written as $(Y_{\nu})_{ij}\equiv c^{\nu}_{ij}\epsilon^{n^{\nu}_{ij}}$ and yields $Y_{\nu}=U_{\nu}\hat{Y}_{\nu}W_{\nu}^{\dagger}~\implies~\hat{m}_{\nu}=\frac{v}{\sqrt{2}}\hat{Y}_{\nu},$ where $\hat{Y}_{\nu}$ is the diagonal Yukawa matrix, $U_{\nu}$ and $W_{\nu}$ are unitary matrices, $\hat{m}_{\nu}$ is the diagonal mass matrix, and the PMNS matrix can be written as $V\equiv U^{\dagger}_{\ell}U_{\nu}$.
@@
Majorana – The Weinberg operator [Weinberg:1979sa]
$\mathcal{L}_{W}\!\!\supset\!-\frac{c_{ij}^{W}\epsilon^{n^{W}_{ij}}}{\Lambda_{W}}(L_{i}H)(L_{j}H),$ (3)
@@
Seesaw – We consider a type-I seesaw scenario [Minkowski:1977sc, Yanagida:1980xy, Gell-Mann:1979vob, Mohapatra:1979ia] where flavor breaking and lepton violation are governed by the FN mechanism. In this case, neutrinos receive both Dirac and Majorana mass terms:
@@
where $c^{\nu}$ and $c^{M}$ are matrices of order-one coefficients, $M$ is the Majorana mass scale of the RH neutrinos, and $n^{\nu}_{ij}\equiv|X_{L_{i}}+X_{N_{j}}|,n^{M}_{ij}\equiv|X_{N_{i}}+X_{N_{j}}|$. In the seesaw limit, the Dirac contribution is small and the diagonal neutrino mass matrix is
$\hskip 11.38092pt\hat{m}_{\nu}\approx\frac{v^{2}}{2M}U_{\nu}^{T}\!\left(c^{\nu}\epsilon^{n^{\nu}}\right)\!\!\left(c^{M}\epsilon^{n^{M}}\right)^{-1}\!\left(c^{\nu}\epsilon^{n^{\nu}}\right)^{T}\!U_{\nu},$~~~~~ (6)
@@
Note that most explicit models of neutrino mass generation can be accommodated within these scenarios [Magg:1980ut, Schechter:1980gr, Lazarides:1980nt, Mohapatra:1980yp, Wetterich:1981bx] (see Appendix C.3).
@@ III Methods
Identifying Realistic Natural FN Textures – We begin by identifying viable FN textures that adequately reproduce the observed lepton masses and mixing parameters for a common value of $\epsilon$. To ensure that, for a given texture, all hierarchical structure arises only from $\epsilon$, we demand that all other free parameters are order-one numbers. In our analysis, these free parameters are encoded in the Lagrangian coefficients $c^{\ell}_{ij},c^{\nu}_{ij},c^{W}_{ij},$ and $c^{M}_{ij}$ from Eqs. (1), (2), (3), and (5), respectively. We scan over a wide range of possible FN charges and determine the fraction of random ${\cal O}(1)$ values in these coefficients that reproduce all observed masses and mixings within some tolerance.
@@
Specifically, we adopt and extend the Bayesian-inspired method of Ref. [Cornella:2023zme] to the lepton sector – see Appendix A for details. For each mass generation mechanism, we consider all textures with charges $|X|\leq 7$, and, for Dirac RH neutrinos, $|X|\leq 9$ as required to obtain viable models. For each texture, we generate random coefficient matrices $c_{ij}$, with each $\log_{10}c_{ij}$ sampled from a normal distribution centered on zero with standard deviation $\sigma=0.3$, and phases sampled uniformly over $0,2\pi]$. (Alternative choices of reasonable $\mathcal{O}(1)$” priors do not meaningfully affect results.)
For each choice of coefficients, we compute observables $\mathcal{O}=\left\{m_{\ell}~,~\Delta m^{2}_{ij}~,~|V_{ij}|~,~\sum m_{\nu}\right\},$ corresponding to the charged lepton masses, neutrino mass-squared differences, the absolute value of the PMNS matrix elements and the sum of neutrino masses. The fractional deviation of each FN prediction $\mathcal{O}_{\rm FN}$ is $\delta_{\mathcal{O}}\sim\mathcal{O}_{\rm FN}/\mathcal{O}_{\rm exp},$ where $\mathcal{O}_{\rm exp}$ is the experimentally measured value (see Appendix A.1). We maximize over all observables to obtain the overall experimental deviation: $\delta_{\mathrm{max}}\equiv\max_{\mathcal{O}}\,(\delta_{\mathcal{O}}).$ Next, we adjust $\epsilon$ and, where applicable, $\Lambda_{W}$ or $\Lambda$, to minimize $\delta_{\mathrm{max}}$. For each texture, this process is repeated many times for many choices of order-one coefficients. Textures that naturally resemble our world will have $\delta_{\mathrm{max}}\sim\mathcal{O}(1)$ for a large fraction coefficient choices. To compare textures, we define x% $\mathcal{F}_{x}\equiv\text{\% of coeff. choices for which}~\delta_{\mathrm{max}}\leq$ x, allowing for textures to be ranked by $\mathcal{F}_{x}$ for different $x$.
@@
where $c_{ijkl}$ are ${\cal O}(1)$ coefficients and we have defined $n_{ijkl}\equiv|X_{\psi_{i}}-X_{\psi_{j}}+X_{\psi_{k}}-X_{\psi_{l}}|.$ Note that we switched to four-component fermion notation to match SMEFT conventions, but we write $n_{ijkl}$ in terms of the FN charges of the corresponding 2-component fermion fields. For dipole operators, we have
@@
In our analysis, ${\cal O}(1)$ Wilson coefficients are generated in the FN basis at a high scale $\Lambda$ and subsequently rotated into the mass basis using the values of $\epsilon$ and the $U$ and $W$ matrices from Eqs. (4), (6), built using the Lagrangian coefficients appropriate to each scenario. We calculate low-energy observables, including two and three-body CLFV decays of muons and taus, as well as muon-to-electron conversion in nuclei, using the flavio package [Straub:2018kue]. Matching to the relevant energy scales is performed with wilson [Aebischer:2018bkb] (running was found to be a negligible effect). Based on experimental constraints, we derive lower bounds on $\Lambda$. Finally, using these bounds, we estimate the highest possible rates of $e\mu$, $e\tau$, and $\mu\tau$ production at future $e^{+}e^{-}$ and $\mu^{+}\mu^{-}$ colliders. For comparison, we also compute the predictions for a fully anarchic or null texture, where all charges are set to zero, and the Wilson coefficients are assumed to be all $\mathcal{O}(1)$ at a fixed scale. Further details on our methodology can be found in Appendix A.3.